# STI-Bench: Are MLLMs Ready for Precise Spatial-Temporal World Understanding?

### v5

@@ 1 Introduction
The rapid development of Multimodal Large Language Models (MLLMs) [1, 12, 4, 37, 38, 39, 29, 35, 50, 44] has propelled them to the research forefront as a versatile tool to deal with numerous vision and multimodal tasks. Impressive performances have been achieved by MLLMs in general Visual Question Answering tasks [3], which mainly focus on the 2D visual perception and semantic question answering [18, 25, 40, 41, 19, 52, 21].
Beyond 2D visual perception, it has become a prevailing trend to employ MLLMs as an end-to-end solution for Embodied AI [23, 32, 26, 46, 15, 9, 8, 7] and Autonomous Driving [43, 48, 22, 42]. Such tasks require MLLMs to understand the 3D space and time, and then predict optimal manipulation strategies for robotic and vehicular systems. Although many explorations have been conducted, the question remains: Are MLLMs ready for precise spatial-temporal world understanding?
To answer this question, we propose a Spatial-Temporal Intelligence Benchmark (STI-Bench), designed to evaluate MLLMs’ spatial-temporal world understanding capability. We evaluate MLLMs using single video or multiple images as input instead of 3D point clouds. The main reasons are: 1) the majority of state-of-the-art models, e.g., GPT-4o [33] and Gemini [38], can accept images or video as input rather than 3D point clouds; 2) Videos are more frequently used in human’s daily life and they usually contain sufficient information to infer the spatial-temporal environment. STI-Bench contains 300 videos and more than 2,000 QA pairs, covering three major scenarios: Desktop, Indoor, and Outdoor. The videos are sourced from Waymo [36], ScanNet [16], and Omni6DPose [51] respectively, thus encompassing a broad spectrum of real-world environments. As illustrated in Figure 2, we design eight distinct tasks to evaluate models’ ability of static spatial measurement and grounding, and dynamic tasks including speed, acceleration and trajectory estimation.
Through extensive experiments as illustrated in Figure 1, we observe that even the most advanced MLLMs struggle with real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis. Our error analysis reveals three fundamental limitations: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal grounding and integration.
These insights highlight the significant challenges MLLMs face in precisely understanding spatial-temporal information from videos. We believe STI-Bench will serve as an important touchstone that guides the community to distinguish and develop better MLLMs for Embodied AI, Autonomous Driving tasks and beyond.
@@
  • We present STI-Bench, comprising over 300 videos and more than 2,000 tailored questions across outdoor, indoor, and desktop scenarios, providing a systematic quantitative assessment of MLLMs’ spatial-temporal understanding capabilities.
@@ 2.1 Multimodal Large Language Models
MLLMs have achieved groundbreaking performance in visual understanding [1, 4, 12, 38], leveraging large language models (LLMs) [44, 39, 37] and visual encoders. Beyond image-based MLLMs, recent advancements have extended multimodal learning to video understanding. Classical works include models like VideoChat[24], which enable interactive video-based dialogue by integrating multimodal understanding. Subsequent models like Subsequent models like Video-LLaVA[30] enhance visual-language alignment through large-scale vision-language pretraining and fine-tuned adaptation, extending LLaVA[31]’s capability to process video inputs effectively. Recent works, Qwen2.5-VL [44] excels in long-video understanding and temporal localization by incorporating absolute temporal encoding, enabling the model to capture relationships among video frames more effectively. Additionally, its advancement in dynamic resolution modeling allows for seamless adaptation to videos with varying sampling rates, enhancing its versatility in processing diverse video inputs.
@@ 2.2 Spatial Understanding with MLLMs
Video MLLMs have attached great importance on semantic understanding. However, spatial understanding has always been a significant challenge, inspiring recent contribution [9, 11, 14]. This progress represents a significant step toward developing world models and embodied agents. Recent advancements in embodied intelligence have explored integrating large-scale MLLMs into robotic control, enabling better generalization and semantic reasoning. RT-2 [8] introduces a vision-language-action framework that transfers web-scale knowledge to robotic control by representing actions as tokens alongside visual and language data, allowing robots to generalize to novel objects and infer multi-step reasoning tasks. Building on this idea, GR-2 [10] extends generalist robot control across diverse embodiments using a Transformer-based architecture trained on a wide range of robotic tasks, demonstrating adaptability across different platforms. Further refining this approach, $\pi_{0}$ [7] incorporates a flow-matching mechanism to generate continuous, precise action trajectories, enhancing fine-grained manipulation skills. By integrating pretrained MLLMs with an independent action module, $\pi_{0}$ achieves zero-shot task execution and flexible adaptation through fine-tuning. Together, these models highlight the potential of leveraging large-scale learning for robotic control, pushing the boundaries of generalization, task adaptability, and multi-modal reasoning in embodied AI.
@@ 2.3 Video Benchmarks for MLLM
Recently, multiple benchmarks [41, 40, 52, 19] have emerged for comprehensively evaluating MLLMs’ ability of (long) video understanding, especially about visual perception and semantic reasoning in the form of Video Question Answering. LongVideoBench [41] and LVBench [40] focus on the understanding of long videos. Recent published benchmarks like Video-MME [19] and MMBench-Video [18] comprehensively evaluates MLLMs across various video-related tasks. Existing benchmarks primarily focus on high-level semantic understanding, such as entity recognition and event understanding. In addition, they are largely confined to a temporal extension of 2D image understanding, lacking precise 3D spatial and temporal reasoning of physical quantities. Recent works such as VSI-Bench [45], have shed light on a deeper understanding of the natural world by introducing visual-spatial intelligence tasks for MLLMs, where models are required to provide numerical answers in certain scenarios. However, as illustrated in Table 1, the limited inclusion of scenes and spatial-temporal tasks restricts their ability to capture the complexities of the real physical world. In contrast, STI-Bench comprehensively evaluate models’ ability of precise spatial-temporal understanding in tasks of static spatial measurement and physically motion understanding in Desktop, Indoor and Outdoor scenarios.
@@ 3.1 Task Definition
We propose eight tasks in total, each one systematically examining a distinct aspect of MLLMs’ spatial-temporal understanding. We divide these tasks into two main categories: Static Understanding and Dynamic Understanding.
@@
    Dimensional Measurement. Concerns estimates of an object’s geometric size, such as length, width, and height, as well as the distance between objects or between the camera and an object. This requires the ability to transform 2D pixel observations into physical world measurements and accurately perceive depth from monocular inputs.
    Example: ”What is the height of this box?” or ”How close is the camera to the table?”
@@
    Spatial Relation. Focuses on identifying spatial relationships among objects or between the camera and an object, including front and back, left and right, up and below. This task tests models’ ability to understand relative positioning across different reference frames and maintain spatial relationship judgment consistency across varying viewpoints.
    Example: ”Is the chair on the left or right side of the table?” or ”What is the position of the red bag relative to the fur sofa?”
@@
    3D Video Grounding. Given a semantic description such as ”the red backpack on the brown sofa,” the goal is to retrieve the object’s 3D bounding box in the camera coordinate system at a specific point in the video. This requires seamlessly aligning linguistic descriptions with visual features and accurately parameterizing 3D positional information.
    Example: ”Locate the 3D bounding box of the red suitcase near the bed.”
@@
    Displacement and Path Length. Focuses on how far an object or the camera travels between two given time points. This requires tracking consistent reference points across frames and integrating motion information from discrete frames into continuous paths.
    Example: ”How far has the car traveled from 1s to 18s?”
@@
    Speed and Acceleration. Investigates motion parameters by integrating spatial displacement with time intervals. This tests models’ ability to compute spatial derivatives with respect to time and maintain scale consistency across varying distances and perspectives.
    Example: ”What is the average speed of the camera?” or ”How quickly is the ball accelerating?”
@@
    Ego-Centric Orientation. Examines how the camera’s azimuth orientation, parallel to the ground plane, changes over the duration of the video. This requires understanding rotation representations and utilizing fixed scene elements as angular reference points.
    Example: ”How many degrees does the camera’s horizontal orientation shift from the start of the video to its end?”
@@
    Trajectory Description. Describes or infers the camera’s or an object’s motion path throughout the entire video, potentially involving multiple segments of travel and turns. This tests the ability to segment complex trajectories into meaningful components and abstract spatial motion patterns into concise language descriptions.
    Example: ”Summarize the camera trajectory, including distances moved and turns made.”
@@
    Pose Estimation. Given the camera’s initial 3D pose, including position and orientation, estimates its pose at a specified point in the video using only the observed RGB data. This requires visual odometry capabilities and the ability to manage cumulative error in long sequences.
    Example: ”Given the initial pose of the camera, what is the camera’s pose at the requested time?”
Each of these tasks presents unique challenges that collectively evaluate models’ comprehensive spatial-temporal intelligence across different scales, from millimeter-precision desktop manipulation to meter-scale indoor navigation and beyond. Success in these tasks requires not only fundamental 3D spatial reasoning but also physical common sense and the ability to integrate information across different modalities and reference frames over time.
@@ 3.2 Benchmark Construction
During QA pair generation, several issues arose:
  1. 1.
    LLM-generated descriptions could be inaccurate or fail to uniquely identify the target object.
  2. 2.
    Some questions and options remained unreasonable or incorrect, even with detailed guidelines.
  3. 3.
    In certain cases, the video alone did not provide sufficient information. For example, the camera was occluded but lidar data were available.
To address these challenges, we developed a website for multiple rounds of manual filtering and sampling-based review, ensuring high-quality questions. We also randomly shuffled the answer options to enhance evaluation robustness. Ultimately, we curated more than 2,000 high-quality QA pairs from over 300 videos. Details are shown in Figure 3.
@@ 4.1 Settings
We conduct a thorough evaluation of leading MLLMs from diverse model families, focusing on both proprietary and open-source solutions. Specifically, we assess the performance of four proprietary models, GPT-4o[33], Gemini-2.0-Flash[38], Gemini-2.5-Pro[38], and Claude-3.7-Sonnet[2], as well as several representative open-source MLLMs that have undergone specialized video-related training, including Qwen2.5-VL-72B[5], InternVL2.5-78B[13], MiniCPM-V-2.6[47], VideoChat-Flash[27],VideoChat-R1[28] and VideoLLaMA3-7B[49].
Considering the stability of open-source models and the API limitations of proprietary models, we uniformly sample 30 frames from the video for each record and explicitly indicate the sampling FPS for the current video within the prompt. An exception is made for Claude-3.7-Sonnet, for which only 20 frames are sampled due to its API constraints. Our benchmark tasks are presented in a multiple-choice format with five possible answers, hence a random guess baseline yields a 20% accuracy. We measure each model’s accuracy by directly comparing the model’s selected answer with the ground truth, without employing any additional external models or annotations for performance evaluation.
@@ 4.2 Main Results
As shown in Table 2 and Table 3, we present a comprehensive evaluation of various MLLMs on STI-Bench. Overall, Gemini-2.5-Pro achieves the highest average accuracy of 41.4% among all tested models, closely followed by Qwen2.5-VL-72B at 40.7% among open-source models. While these results significantly exceed the random guess baseline (20%), they still highlight substantial room for improvement in spatial-temporal understanding.
When analyzing performance across different scene types (Table 3), we observe varied patterns. Qwen2.5-VL-72B demonstrates notably strong performance in outdoor scenarios (50.6%), while Gemini-2.5-Pro also performs well outdoors (48.7%). In indoor environments, Claude-3.7-Sonnet (38.2%) and Gemini-2.5-Pro (37.1%) lead. For desktop scenes, Gemini-2.5-Pro shows the strongest performance (35.8%), followed by Qwen2.5-VL-72B (33.5%). These variations suggest different model specializations, possibly influenced by training data distribution or architectural differences in handling diverse environmental cues.
Task-specific performance (Table 2) reveals particularly challenging areas. Models generally struggle with tasks requiring precise quantitative estimation. For instance, the best performance on Dimensional Measurement is achieved by Gemini-2.5-Pro (38.7%), and on Displacement & Path Length also by Gemini-2.5-Pro (33.9%). Speed & Acceleration also presents difficulties, with Claude-3.7-Sonnet leading at 38.8%. In contrast, models demonstrate stronger capabilities in Pose Estimation (best: 62.7% by Claude-3.7-Sonnet) and Spatial Relation tasks (best: 53.8% by Gemini-2.5-Pro). Notably, Gemini-2.5-Pro also achieves the highest score in Ego-Centric Orientation (52.5%) among proprietary models, while VideoLLaMA3-7B leads in Trajectory Description (54.6%) among open-source models.
Among open-source models, Qwen2.5-VL-72B stands out with an average accuracy of 40.7%, demonstrating highly competitive performance. However, other open-source models like InternVL2.5-78B (38.5%) also show promise, though smaller models like MiniCPM-V-2.6 (26.9%) still lag behind their larger counterparts and proprietary models.
It is important to emphasize that even the best-performing model, Gemini-2.5-Pro, achieves only 41.4% average accuracy on our benchmark. While this is considerably better than random guessing, it remains far from the reliability required for demanding real-world applications like embodied AI or autonomous driving. These results indicate that current MLLMs, despite their impressive general visual understanding capabilities, still require significant advancements in precise spatial-temporal intelligence for embodied tasks.
@@ 4.3 Experimental Analysis
Given that Gemini-2.5-Pro exhibits strong multi-modal reasoning capabilities, provides detailed thinking processes (as exemplified in Figure 6), and achieves the top overall performance (41.4% average accuracy) among all evaluated models, we select it as a representative for in-depth analysis.
Our analysis of Gemini-2.5-Pro reveals several key characteristics of current state-of-the-art MLLMs’ spatial-temporal understanding capabilities. Overall, the model achieves 41.4% average accuracy. Performance varies across scene types: it performs strongest in outdoor scenarios (48.7%), followed by indoor (37.1%) and desktop environments (35.8%). This performance disparity might suggest that the model’s training data or architecture is better attuned to outdoor and larger-scale understanding, or that such environments typically provide clearer visual cues for spatial reasoning.
When examining task-specific performance for Gemini-2.5-Pro (Table 2), we observe its strongest capabilities in tasks involving relative understanding and object state. It achieves high scores in Spatial Relation (53.8%, the best among all models) and Ego-Centric Orientation (52.5%, the highest among proprietary models), followed by Pose Estimation (50.4%). Interestingly, while tasks requiring precise quantitative estimation generally show lower absolute scores across models, Gemini-2.5-Pro performs very competitively and often leads in these areas: it achieves the top scores in Dimensional Measurement (38.7%) and Displacement & Path Length (33.9%). Its score for Speed & Acceleration is 33.1%. This indicates that while the absolute accuracy in precise estimation tasks remains a challenge, Gemini-2.5-Pro demonstrates comparatively strong capabilities in these demanding domains relative to other models.
By leveraging the model’s reasoning process and uniformly sampling approximately 200 error records across each task type and scenario, we categorize its errors into three representative patterns. Figure 7 shows the distribution of error categories.
@@
The model faces significant challenges in accurately estimating static spatial properties and relationships from visual inputs. These difficulties manifest when estimating object dimensions (size), distances between objects or between camera and objects, and absolute 3D coordinates/bounding boxes at specific time points. These errors stem from a lack of clear visual size references, difficulty distinguishing between numerically close options, and the inherent challenges of inferring metric scale from 2D pixels and estimating depth with monocular cameras. Such limitations directly impact performance in dimensional measurement, spatial relation, and 3D video grounding tasks.
@@
The model struggles to perceive, track, and interpret cross-frame information that changes over time, such as motion and its dynamics. This results in erroneous calculations or descriptions of displacement, path length, speed, acceleration, directional changes (ego-centric or object pose), and overall trajectory shapes. The model particularly struggles with relative motion (distinguishing object motion from camera motion), a problem exacerbated by sparse temporal sampling. These difficulties arise from challenges in integrating information across frames, lack of internal models for physics/kinematics, inability to separate ego-motion from object motion, and information loss due to sparse sampling. These issues manifest in tasks involving displacement and path length, speed and acceleration, ego-centric orientation, trajectory description, and pose estimation (as it changes over time).
@@
The model fails to properly connect textual queries/instructions with relevant spatial-temporal visual content, or to integrate provided non-visual data (such as initial poses) with visual information. This includes misinterpreting temporal constraints (like ”from 1s to 18s,” ”at the end,” ”the moment of last co-occurrence”), failing to correctly utilize provided initial conditions (e.g., initial camera pose in pose estimation tasks), and incorrectly associating structured data (coordinates, timestamps) with visual elements. These errors stem from deficiencies in parsing structured/natural language instructions and difficulty integrating information from different modalities (text prompts, initial state data, video frames) into a unified reasoning process. This affects all tasks that rely on specific instructions or initial data.

### v6

@@ 1 Introduction
Recent advances in Multimodal Large Language Models (MLLMs) [1, 12, 4, 37, 38, 39, 29, 35, 50, 44] have positioned them as powerful tools for numerous vision and multimodal tasks. MLLMs achieve impressive performance in general Visual Question Answering tasks [3], which mainly focus on 2D visual perception and semantic question answering [18, 25, 40, 41, 19, 52, 21]. Beyond 2D visual perception, MLLMs are increasingly employed as end-to-end solutions for Embodied AI [23, 32, 26, 46, 15, 9, 8, 7] and Autonomous Driving [43, 48, 22, 42]. Such tasks require MLLMs to understand 3D space and time, and predict optimal manipulation strategies for robotic and vehicular systems. Despite these explorations, the question remains: Are MLLMs ready for precise spatial-temporal world understanding?
To answer this question, we propose a Spatial-Temporal Intelligence Benchmark (STI-Bench), designed to evaluate MLLMs’ spatial-temporal world understanding capability. We evaluate MLLMs using single video or multiple images as input instead of 3D point clouds. The main reasons are: 1) the majority of state-of-the-art models, e.g., GPT-4o [33] and Gemini [38], accept images or video as input rather than 3D point clouds; 2) Videos are more frequently used in daily life and contain sufficient information to infer the spatial-temporal environment.
STI-Bench contains 300 videos and over 2,000 QA pairs, covering three major scenarios: Desktop, Indoor, and Outdoor. The videos are sourced from Waymo [36], ScanNet [16], and Omni6DPose [51] respectively, encompassing a broad spectrum of real-world environments. As illustrated in Figure 2, we design eight distinct tasks to evaluate models’ ability in static spatial measurement and grounding, and dynamic tasks including speed, acceleration and trajectory estimation.
Through extensive experiments as illustrated in Figure 1, we observe that even the most advanced MLLMs struggle with real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis. Our error analysis reveals three fundamental limitations: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal grounding and integration. These insights highlight the significant challenges MLLMs face in precisely understanding spatial-temporal information from videos. We believe STI-Bench will serve as an important touchstone that guides the community to develop better MLLMs for Embodied AI, Autonomous Driving tasks and beyond.
@@
  • We present STI-Bench, comprising over 300 videos and over 2,000 tailored questions across outdoor, indoor, and desktop scenarios, providing a systematic quantitative assessment of MLLMs’ spatial-temporal understanding capabilities.
@@ 2.1 Multimodal Large Language Models
MLLMs have achieved groundbreaking performance in visual understanding [1, 4, 12, 38], leveraging large language models (LLMs) [44, 39, 37] and visual encoders. Recent advancements have extended multimodal learning to video understanding. Classical works include VideoChat[24], which enables interactive video-based dialogue by integrating multimodal understanding. Subsequent models like Video-LLaVA[30] enhance visual-language alignment through large-scale vision-language pretraining, extending LLaVA[31]’s capability to process video inputs effectively. Recent works like Qwen2.5-VL [44] excel in long-video understanding and temporal localization by incorporating absolute temporal encoding, enabling models to capture relationships among video frames more effectively.
@@ 2.2 Spatial Understanding with MLLMs
Video MLLMs have focused heavily on semantic understanding. However, spatial understanding remains a significant challenge, inspiring recent contributions [9, 11, 14]. This progress represents a significant step toward developing world models and embodied agents. Recent advancements in embodied intelligence have explored integrating large-scale MLLMs into robotic control, enabling better generalization and semantic reasoning. RT-2 [8] introduces a vision-language-action framework that transfers web-scale knowledge to robotic control by representing actions as tokens alongside visual and language data. Building on this idea, GR-2 [10] extends generalist robot control across diverse embodiments using a Transformer-based architecture. Further refining this approach, $\pi_{0}$ [7] incorporates a flow-matching mechanism to generate continuous, precise action trajectories, enhancing fine-grained manipulation skills.
@@ 2.3 Video Benchmarks for MLLM
Recently, multiple benchmarks [41, 40, 52, 19] have emerged for comprehensively evaluating MLLMs’ ability in (long) video understanding, especially visual perception and semantic reasoning through Video Question Answering. LongVideoBench [41] and LVBench [40] focus on long video understanding, while Video-MME [19] and MMBench-Video [18] comprehensively evaluate MLLMs across various video-related tasks. Existing benchmarks primarily focus on high-level semantic understanding, such as entity recognition and event understanding, largely confined to a temporal extension of 2D image understanding, lacking precise 3D spatial and temporal reasoning of physical quantities. Recent works such as VSI-Bench [45] have introduced visual-spatial intelligence tasks for MLLMs, where models are required to provide numerical answers in certain scenarios. However, as illustrated in Table 1, the limited inclusion of scenes and spatial-temporal tasks restricts their ability to capture the complexities of the real physical world. In contrast, STI-Bench comprehensively evaluates models’ ability in precise spatial-temporal understanding through tasks of static spatial measurement and physical motion understanding in Desktop, Indoor and Outdoor scenarios.
@@ 3.1 Task Definition
We propose eight tasks examining distinct aspects of MLLMs’ spatial-temporal understanding, divided into Static Understanding and Dynamic Understanding.
@@
    Dimensional Measurement. Estimates object geometric size and distances, requiring transformation from 2D pixel observations to physical measurements.
    Example: "What is the height of this box?"
@@
    Spatial Relation. Identifies spatial relationships among objects or between camera and objects, testing relative positioning understanding across viewpoints.
    Example: "Is the chair on the left or right of the table?"
@@
    3D Video Grounding. Retrieves object’s 3D bounding box given semantic descriptions, requiring alignment of linguistic and visual features.
    Example: "Locate the 3D bounding box of the red suitcase near the bed."
@@
    Displacement and Path Length. Measures travel distance between time points, requiring motion tracking across frames.
    Example: "How far has the car traveled from 1s to 18s?"
@@
    Speed and Acceleration. Computes motion parameters by integrating spatial displacement with time intervals.
    Example: "What is the average speed of the camera?"
@@
    Ego-Centric Orientation. Examines camera’s azimuth orientation changes, requiring understanding of rotation representations.
    Example: "How many degrees does the camera’s horizontal orientation shift?"
@@
    Trajectory Description. Describes motion paths throughout the video, testing ability to abstract spatial motion into language.
    Example: "Summarize the camera trajectory, including distances moved and turns made."
@@
    Pose Estimation. Estimates camera pose at specified time points using RGB data, requiring visual odometry capabilities.
    Example: "Given the initial pose, what is the camera’s pose at the requested time?"
These tasks collectively evaluate comprehensive spatial-temporal intelligence across different scales, requiring fundamental 3D spatial reasoning, physical common sense, and cross-modal information integration over time.
@@ 3.2 Benchmark Construction
During QA pair generation, several issues arose: LLM-generated descriptions could be inaccurate or fail to uniquely identify the target object; some questions and options remained unreasonable or incorrect, even with detailed guidelines; and in certain cases, the video alone did not provide sufficient information (e.g., when the camera was occluded but lidar data were available). To address these challenges, we developed a website for multiple rounds of manual filtering and sampling-based review, ensuring high-quality questions. We also randomly shuffled the answer options to enhance evaluation robustness. Ultimately, we curated more than 2,000 high-quality QA pairs from over 300 videos. Details are shown in Figure 3.
@@ 4.1 Settings
We conduct a thorough evaluation of leading MLLMs from diverse model families, focusing on both proprietary and open-source solutions. Specifically, we assess four proprietary models: GPT-4o[33], Gemini-2.0-Flash[38], Gemini-2.5-Pro[38], and Claude-3.7-Sonnet[2], as well as several representative open-source MLLMs that have undergone specialized video-related training, including Qwen2.5-VL-72B[5], InternVL2.5-78B[13], MiniCPM-V-2.6[47], VideoChat-Flash[27], VideoChat-R1[28] and VideoLLaMA3-7B[49].
Considering the stability of open-source models and the API limitations of proprietary models, we uniformly sample 30 frames from the video for each record and explicitly indicate the sampling FPS within the prompt. An exception is made for Claude-3.7-Sonnet, for which only 20 frames are sampled due to its API constraints. Our benchmark tasks are presented in a multiple-choice format with five possible answers, hence a random guess baseline yields 20% accuracy. We measure each model’s accuracy by directly comparing the model’s selected answer with the ground truth.
@@ 4.2 Main Results
As shown in Table 2 and Table 3, we present a comprehensive evaluation of various MLLMs on STI-Bench. Gemini-2.5-Pro achieves the highest average accuracy of 41.4%, closely followed by Qwen2.5-VL-72B at 40.7% among open-source models. While these results significantly exceed the random guess baseline (20%), they highlight substantial room for improvement in spatial-temporal understanding.
Performance across different scene types shows varied patterns. Qwen2.5-VL-72B demonstrates strong performance in outdoor scenarios (50.6%), while Gemini-2.5-Pro performs well outdoors (48.7%) and leads in desktop scenes (35.8%). In indoor environments, Claude-3.7-Sonnet (38.2%) and Gemini-2.5-Pro (37.1%) lead. These variations suggest different model specializations, possibly influenced by training data distribution.
Task-specific performance reveals particularly challenging areas. Models generally struggle with precise quantitative estimation tasks. The best performance on Dimensional Measurement is achieved by Gemini-2.5-Pro (38.7%), and on Displacement & Path Length also by Gemini-2.5-Pro (33.9%). In contrast, models demonstrate stronger capabilities in Pose Estimation (best: 62.7% by Claude-3.7-Sonnet) and Spatial Relation tasks (best: 53.8% by Gemini-2.5-Pro).
Among open-source models, Qwen2.5-VL-72B stands out with 40.7% average accuracy, demonstrating highly competitive performance. Other open-source models like InternVL2.5-78B (38.5%) also show promise, though smaller models like MiniCPM-V-2.6 (26.9%) lag behind their larger counterparts.
Even the best-performing model, Gemini-2.5-Pro, achieves only 41.4% average accuracy on our benchmark. These results indicate that current MLLMs, despite impressive general visual understanding capabilities, still require significant advancements in precise spatial-temporal intelligence for embodied tasks.
@@ 4.3 Experimental Analysis
Given that Gemini-2.5-Pro exhibits strong multi-modal reasoning capabilities and achieves the top overall performance (41.4% average accuracy), we select it as a representative for in-depth analysis.
Performance varies across scene types: it performs strongest in outdoor scenarios (48.7%), followed by indoor (37.1%) and desktop environments (35.8%). This disparity might suggest that the model’s training data is better attuned to outdoor and larger-scale understanding.
Examining task-specific performance, we observe its strongest capabilities in tasks involving relative understanding and object state. It achieves high scores in Spatial Relation (53.8%) and Ego-Centric Orientation (52.5%), followed by Pose Estimation (50.4%). While tasks requiring precise quantitative estimation show lower absolute scores, Gemini-2.5-Pro performs competitively and often leads in these areas, achieving top scores in Dimensional Measurement (38.7%) and Displacement & Path Length (33.9%).
By leveraging the model’s reasoning process and uniformly sampling approximately 200 error records across each task type and scenario, we categorize errors into three representative patterns. Figure 7 shows the distribution of error categories.
@@
The model faces significant challenges in accurately estimating static spatial properties and relationships from visual inputs. These difficulties manifest when estimating object dimensions, distances between objects or between camera and objects, and absolute 3D coordinates at specific time points. These errors stem from a lack of clear visual size references, difficulty distinguishing between numerically close options, and the inherent challenges of inferring metric scale from 2D pixels and estimating depth with monocular cameras. Such limitations directly impact performance in dimensional measurement, spatial relation, and 3D video grounding tasks.
@@
The model struggles to perceive, track, and interpret cross-frame information that changes over time, such as motion and its dynamics. This results in erroneous calculations or descriptions of displacement, path length, speed, acceleration, directional changes, and overall trajectory shapes. The model particularly struggles with relative motion (distinguishing object motion from camera motion), a problem exacerbated by sparse temporal sampling. These difficulties arise from challenges in integrating information across frames, lack of internal models for physics/kinematics, inability to separate ego-motion from object motion, and information loss due to sparse sampling. These issues manifest in tasks involving displacement and path length, speed and acceleration, ego-centric orientation, trajectory description, and pose estimation.
@@
The model fails to properly connect textual queries/instructions with relevant spatial-temporal visual content, or to integrate provided non-visual data (such as initial poses) with visual information. This includes misinterpreting temporal constraints (like "from 1s to 18s," "at the end," "the moment of last co-occurrence"), failing to correctly utilize provided initial conditions (e.g., initial camera pose in pose estimation tasks), and incorrectly associating structured data (coordinates, timestamps) with visual elements. These errors stem from deficiencies in parsing structured/natural language instructions and difficulty integrating information from different modalities into a unified reasoning process. This affects all tasks that rely on specific instructions or initial data.